# Low-Rank Interconnected Adaptation across Layers

### v2

@@ 4 Experiments
We validate the effectiveness of Lily across different domains, model sizes (from ViT to LLM), and architectures (Transformers, Mamba), demonstrating its general strong adaptation capability. Concurrently, we conduct a comprehensive analysis of Lily’s intrinsic mechanisms, providing a thorough understanding of Lily. All experiments are conducted on a single RTX 4090 GPU. Additionally, multiple analysis are provided in Appendix C, D, E, F, G, H, I and J.
@@ 4.1 Common Sense Reasoning
Table 1: Commonsense reasoning results for Falcon-Mamba-7B across eight tasks. Bold represents the highest performance for each dataset utilizing PEFT methods.
@@
Model PEFT BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg.
ChatGPT - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
Falcon-Mamba-7B LoRA 6.5 30.5 40.6 14.9 56.4 42.2 31.8 38.4 32.7
Lily (Δ + $\Delta\text{ + in}$) 44.9 66.8 65.0 10.5 57.1 78.7 64.6 68.2 57.0
Lily (in) 60.2 61.0 67.3 12.9 61.5 80.0 67.5 65.8 59.5
@@
Table 2: Commonsense reasoning results for LLaMA3-8B across eight tasks. represents results taken from Liu et al. (2024). Bold denotes the highest performance scores for each dataset among different PEFT methods.
@@
Model PEFT BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg.
ChatGPT - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
LLaMA3-8B LoRA 70.8 85.2 79.9 91.7 84.3 84.2 71.2 79.0 80.8
PiSSA 67.1 81.1 77.2 83.6 78.9 77.7 63.2 74.6 75.4
MiLoRA 68.8 86.7 77.2 92.9 85.6 86.8 75.5 81.8 81.9
Lily 72.9 85.6 77.8 92.7 83.3 89.7 77.6 82.8 82.8
@@
Implementation We evaluate Lily on commonsense reasoning with LLMs. Regarding the implementation, we utilize LLaMA3-8B (AI@Meta, 2024) and Falcon-Mamba-7B (Zuo et al., 2024) as backbones. LLaMA3 is a near-SOTA open-source large language model, while Falcon-Mamba is the latest and only open-source large language model based on the Mamba architecture. Using these models allows us to validate the effectiveness of Lily for fine-tuning LLMs and whether this effectiveness can be transferred to architectures beyond Transformers (Mamba, in this case). We fine-tune these models on Commonsense170K (Hu et al., 2023) and evaluate the adaptation results on eight multiple-choice problem tasks, including BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e, ARC-c (Clark et al., 2018), and OBQA (Mihaylov et al., 2018). The compared methods are LoRA for Falcon-Mamba and LoRA (Hu et al., 2021), PiSSA (Meng et al., 2024), and MiLoRA (Wang et al., 2024) for LLaMA3. We only compare LoRA for Falcon-Mamba because tailored PEFT methods for Mamba-based LLMs have not yet been proposed, which is beyond the scope of this paper. Detailed hyper-parameter settings and datsets information are reported in Appendix B.1.1 and Appendix B.2.1.
@@
Results We report the accuracy in the Table 2 and Table 1. Based on the results, it is evident that Lily performs the best out of the compared PEFT methods. Lily surpasses LoRA by a significant margin on Falcon-Mamba, and on LLaMA3, it outperforms LoRA and MiLoRA. This indicates Lily’s superior adaptation capability and parameter efficiency dealing with commonsense reasoning tasks. Additionally, while the performance on Falcon-Mamba is notably lower than the baseline and LLaMA3, we believe this is due to the model’s limitations rather than Lily’s, as Lily still significantly outperforms LoRA on Falcon-Mamba and demonstrates great performance on LLaMA3. This sheds light on the current state of Mamba-based LLMs, showing that they generally have inferior performance compared to Transformer-based LLMs like ChatGPT and LLaMA on many tasks.
@@ 4.2 Natural Language Understanding
Table 3: Various fine-tuning methods applied to RoBERTa Base are evaluated on 6 datasets from the GLUE benchmark. We present the Matthew’s correlation coefficient (MCC) for CoLA, Pearson correlation coefficient (PCC) for STS-B, and accuracy (Acc.) for the remaining tasks. The highest performance for each dataset is highlighted in bold, with all metrics favoring higher values across the 6 datasets.
@@
Implementation We evaluate Lily on natural language understanding (NLU) tasks. For implementation, we use RoBERTa Base (Liu et al., 2019) as the backbone and fine-tune it on tasks from GLUE benchmark (General Language Understanding Evaluation (Wang et al., 2018)), consisting of multiple NLU tasks including single-sentence classification tasks, similarity and paraphrase tasks and natural language inference tasks. We compare Lily against several competitive PEFT methods, including BitFit (Zaken et al., 2021), Adapter-Tuning (Rücklé et al., 2020) (Houlsby et al., 2019a) (Lin et al., 2020) (Pfeiffer et al., 2020a), LoRA (Hu et al., 2021), DyLoRA (Valipour et al., 2022) and AdaLoRA (Zhang et al., 2023). Additionally, we utilize full fine-tuning (FFT) as the baseline. Specific hyper-parameters and datasets information are provide in Appendix B.1.2 and B.2.2.
Results The results are shown in Table. 3, from which we can clearly observe that Lily surpass all of the compared PEFT methods by a significant margin, demonstrating its capability of tackling NLU tasks. Among the 6 given tasks, Lily surpasses FFT in 4 of them, showcasing its strong approximation ability with compatible parameter-efficiency.
Figure 2: Results of subject-driven generation. Lily’s results align better with prompts, featuring more accurate color, environment, and shape.
@@ 4.3 Subject-driven Image Generation
Implementation We conduct experiments on fine-tuning text-to-image diffusion models for the subject-driven generation task (Ruiz et al., 2023). For backbone, we use SDXL and we fine-tune it using LoRA and Lily. We first fine-tune the model with images associated with text prompts (e.g., A photo of a [v] duck toy), in which a unique identifier is provided. After that, text prompts containing the identifier could be used to generate customized images.
Results The results are presented in Fig. 2 following the format in Gao et al. (2024b) and Wu et al. (2024), from which we can observe that images generated by Lily generally align better with the text prompts. For instance, when asked to generate a duck toy floating on top of water, Lily’s image accurately depicts the designated environment, whereas LoRA’s does not. Additionally, when asked to generate a wolf plushie in snow, Lily precisely depicts the snow around the wolf, while LoRA fails to do so. These observations demonstrate Lily’s excellent ability in the domain of text-to-image generation with more expressive adaptation. More generated results are in Appendix I.
@@ 4.4 Visual Adaptation Benchmark
Table 5: Full results of Lily on Vim-S pre-trained on ImageNet-1K for the VTAB-1K benchmark, with averages calculated within each group. * denotes linear probing results from Tu et al. (2023b). For fair comparison, we also use ViT-B pre-trained on ImageNet-1K. Bold indicates best performance among Vim-based PEFT methods.
@@
Implementation We assess Lily on the Visual Task Adaptation Benchmark (VTAB-1K Zhai et al. (2019)), a suite of 19 visual tasks spanning diverse domains and semantics, to test its general visual adaptation capability. Tasks are categorized into Natural, Specialized, and Structured, all formulated as classification problems for consistent model evaluation. We conduct two sets of experiments: one focusing on the adaptation effectiveness on Vision Transformer (ViT (Dosovitskiy et al., 2020)) and the other on Vision Mamba (Vim (Zhu et al., 2024)), demonstrating Lily’s architecture-agnostic capabilities. For ViT, we use ViT-B pre-trained on ImageNet-21K (Deng et al., 2009), and for Vim, Vim-s pre-trained on ImageNet-1K. To fairly compare ViT and Vim architectures, we implement LoRA (Hu et al., 2021) and AdaptFormer (Chen et al., 2022) on ViT-B pre-trained on ImageNet-1K. In ViT experiments, we compare Lily with LoRA, AdaptFormer, FourierFT (Gao et al., 2024b), and MoRA (Jiang et al., 2024); in Vim experiments, we focus on contrasting architectures difference, therefore only using LoRA as the baseline. All experiments include FFT and linear probing as baselines. For Vim, we implement two versions: Lily-S (Small) and Lily-L (Large) of Lily, with different hyperparameter settings to either reduce the parameter count (Lily-S) or maximize performance (Lily-L). For Lily on ViT, the reported results are obtained from adapting both the self attention and the MLP module in Transformer. For the performance w.r.t the fine-tuned module, we conduct additional experiments in Appendix D. Detailed experimental settings and datasets information are provide in Appendix B.1.3 and B.2.3.
Results Results are shown in Table 4 and Table 5. For ViT, Lily significantly outperforms all compared PEFT methods with improved parameter efficiency. For Vim, results on ViT generally surpass those on Vim. For instance, LoRA on ViT performs better than LoRA on Vim. We argue that this is due to differences in architecture designs and general model sizes. However, Lily’s strong adaptation performance allows it to match or exceed PEFT methods on ViT and significantly outperform LoRA on Vim (Lily-S and Lily-L surpass LoRA by a significant margin). This demonstrates Lily’s architecture-agnostic capability, highlighting its potential across various model architectures. In general, Lily has achieved great visual adaptation capability with an advantage of being architecture-agnostic and enjoying excellent parameter-efficiency.
@@ 4.5 Understanding Lily
From a Feature Merging Perspective. Apart from being a better gradient projector than LoRA, Lily also enables comprehensive information access across layers. Lily enables access to information or features from all other layers when adapting a target module at a specific layer thanks to the inter-connectivity of the adapters. We aim to understand how Lily achieves this comprehensive information access from the perspective of visual tasks as shown in Fig. 3. We can observe that, in Lily, the distinctness of the attention maps between layers is not as pronounced as in LoRA. This validates Lily’s ability to enable all-level information access, since adaptation at each layer takes into account features from other layers. Additionally, we specifically visualize the actual feature differences between different layers in Fig. 4. We observe that Lily has more points with low feature differences (blue color) than LoRA, indicating that the distinctness of features between layers in Lily is generally lower than in LoRA. This further demonstrates Lily’s ability to enable comprehensive information access. Although we enable all-level information access, what prevents the features from becoming completely identical is the selectivity introduced by Lily, which we specify in the following section.
Figure 5: Impact of attention granularity (i.e., the choice of how many LPs and HPs) on the performance. We choose 12 out of 19 tasks from VTAB-1K for a comprehensive understanding.
@@
Figure 6: Visualization of accumulated assigned weight for HP experts by a router across various layers. Example here uses layer of index 2, 13 and 22 to represent shallow, middle and deep layers.
@@
What’s the Influence of Attention Granularity? The number of experts in the model-wide HP module can be freely set, and the number of LPs can also be flexibly set by sharing across the same level of layers introduced in Appendix A.1. Therefore, we analyze the impact of these choices on performance. We denote the number of LP experts and HP experts as ne_1 and ne_2, respectively. For simplicity, we make them identical in the experiments, denoted as ne. We refer to the number of layers each expert attends to as attention granularity. As the value of ne increases, the attention granularity becomes finer. As shown in Fig. 5, the results from the VTAB-1K benchmark indicate different patterns. For instance, on the DTD dataset, the best performance is achieved when ne is 4, while on sNORB-Azim, performance increases with the increase in ne. Increasing ne leads to more parameters and finer attention granularity. However, finer attention granularity does not necessarily lead to better overall performance. For example, on Resisc45, DTD, Cifar100, sNORB-Ele, dsPr-LoC, Flowers102, and EuroSAT, the negative impact of increasingly finer attention granularity eventually outweighs the benefits of increased parameters, leading to a decrease in overall performance. In other tasks, different patterns may occur because the positive effect of attention granularity on performance is consistently strong, or its negative effect is not enough to offset the benefits of increased parameters, resulting in a generally increasing performance with ne. This phenomenon provides an important insight: for most tasks, simply increasing parameters may not lead to better performance. Instead, only when attention granularity and the number of parameters reach a good tradeoff can we achieve the best performance.

### v3

@@ 4 Experiments
We validate the effectiveness of Lily across different domains, model sizes (from ViT to LLM), and architectures (Transformers, Mamba), demonstrating its generally strong adaptation capability. Concurrently, we conduct a comprehensive analysis of Lily’s intrinsic mechanisms, providing a thorough understanding of how it works. All ranks for Lily are selected from 8, 16, 32, ensuring that the total parameter count does not exceed that of the baselines. All experiments are conducted on a single RTX 4090 GPU. Additionally, multiple analyses are provided in Appendix C, D, E, F, G, H, I, and J.
@@ 4.1 Common Sense Reasoning
Table 1: Commonsense reasoning results for Falcon-Mamba-7B across eight tasks. Bold represents the highest performance for each dataset utilizing PEFT methods. “$\Delta$” and “in” refer to adaptations of Mamba’s delta_proj and in_proj parameters, respectively.
@@
Model PEFT Params BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg.
ChatGPT - - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
Falcon-Mamba-7B LoRA 3.7M 6.5 30.5 40.6 14.9 56.4 42.2 31.8 38.4 32.7
Lily (Δ + $\Delta\text{ + in}$) 3.7M 44.9 66.8 65.0 10.5 57.1 78.7 64.6 68.2 57.0
Lily (in) 3.3M 60.2 61.0 67.3 12.9 61.5 80.0 67.5 65.8 59.5
@@
Table 2: Commonsense reasoning results for LLaMA3-8B across eight tasks. represents results taken from Liu et al. (2024) and (Wang et al., 2024). Bold denotes the highest performance scores for each dataset among different PEFT methods.
@@
Model PEFT Params BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg.
ChatGPT - - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
LLaMA3-8B LoRA 56M 70.8 85.2 79.9 91.7 84.3 84.2 71.2 79.0 80.8
PiSSA 83.8M 67.1 81.1 77.2 83.6 78.9 77.7 63.2 74.6 75.4
MiLoRA 56.6M 68.8 86.7 77.2 92.9 85.6 86.8 75.5 81.8 81.9
Lily 1.2M 72.9 85.6 77.8 92.7 83.3 89.7 77.6 82.8 82.8
@@
Implementation: We evaluate Lily on commonsense reasoning with LLMs. For the implementation, we utilize LLaMA3-8B (AI@Meta, 2024) and Falcon-Mamba-7B (Zuo et al., 2024) as backbones. LLaMA3 is a near-SOTA open-source large language model, while Falcon-Mamba is an open-sourced large language model based on the Mamba architecture. Using these models allows us to validate the effectiveness of Lily for fine-tuning LLMs and assess whether this effectiveness can be transferred to architectures beyond Transformers (Mamba, in this case). We fine-tune these models on Commonsense170K (Hu et al., 2023) and evaluate the adaptation results on eight multiple-choice problem tasks, including BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e, ARC-c (Clark et al., 2018), and OBQA (Mihaylov et al., 2018). The compared methods are LoRA for Falcon-Mamba and LoRA (Hu et al., 2021), PiSSA (Meng et al., 2024), and MiLoRA (Wang et al., 2024) for LLaMA3. We only compare LoRA for Falcon-Mamba because tailored PEFT methods for Mamba-based LLMs have not yet been proposed, which is beyond the scope of this paper. Detailed hyper-parameter settings and dataset information are reported in Appendix B.1.1 and Appendix B.2.1.
@@
Results We report the accuracy in Tables 2 and 1. Based on these results, it can be observed that Lily outperforms the other compared PEFT methods with a smaller parameters budget. Specifically, Lily surpasses LoRA by a significant margin on Falcon-Mamba and, on LLaMA3, outperforms both LoRA and MiLoRA. This demonstrates Lily’s superior adaptation capability and parameter efficiency in handling commonsense reasoning tasks. Additionally, although performance on Falcon-Mamba is notably lower than that of the baseline and LLaMA3, we believe this discrepancy stems from the inherent limitations of the model rather than any deficiency in Lily, as Lily still significantly outperforms LoRA on Falcon-Mamba while demonstrating robust performance on LLaMA3. These findings also highlight that current state of Mamba-based LLMs generally exhibits inferior performance compared to Transformer-based LLMs such as ChatGPT and LLaMA on many tasks.
@@ 4.2 Natural Language Understanding
Table 3: Various fine-tuning methods applied to RoBERTa Base and RoBERTa Large are evaluated on six datasets from the GLUE benchmark. We present the Matthews correlation coefficient (MCC) for CoLA, the Pearson correlation coefficient (PCC) for STS-B, and accuracy (Acc.) for the remaining tasks. The highest performance for each dataset is highlighted in bold, with all metrics favoring higher values across the six datasets.
@@
$\rm{RoB_{large}}$(FF) 356M 96.4 90.9 68 94.7 86.6 92.4 88.2
$\rm{RoB_{large}}$($\text{Adpt}^{\text{H}}$) 0.8M 96.3 87.7 66.3 94.7 72.9 91.5 84.9
$\rm{RoB_{large}}$(LoRA) 0.8M 96.2 90.2 68.2 94.8 85.2 92.3 87.8
$\rm{RoB_{large}}$(Lily) 0.5M 95.6 90.9 68.4 94.8 88.4 91.9 88.4
Implementation We evaluate Lily on natural language understanding (NLU) tasks. For the implementation, we use RoBERTa Base (Liu et al., 2019) and RoBERTa Large as the backbones and fine-tune them on tasks from the GLUE benchmark (General Language Understanding Evaluation (Wang et al., 2018)), which consists of multiple NLU tasks, including single-sentence classification, similarity and paraphrase, and natural language inference tasks. We compare Lily against several competitive PEFT methods, including BitFit (Zaken et al., 2021), Adapter-Tuning (Rücklé et al., 2020; Houlsby et al., 2019b; Lin et al., 2020; Pfeiffer et al., 2020b), LoRA (Hu et al., 2021), DyLoRA (Valipour et al., 2022), FLoRA (Hao et al., 2024), and AdaLoRA (Zhang et al., 2023). Additionally, we utilize full fine-tuning (FFT) as the baseline. Specific hyperparameters and dataset information are provided in Appendix B.1.2 and B.2.2.
Results The results are shown in Table 3. From the table, we can clearly observe that Lily surpasses all the compared PEFT methods by a significant margin, demonstrating its ability to tackle NLU tasks. Among the six tasks, Lily surpasses FFT on four of them when using RoBERTa Base and RoBERTa Large, showcasing its strong approximation ability and high parameter efficiency.
Figure 2: Qualitative results of subject-driven generation. Lily’s results align better with prompts, featuring more accurate color, environment, and shape.
@@ 4.3 Subject-driven Image Generation
Implementation We conduct experiments on fine-tuning text-to-image diffusion models for the subject-driven generation task. As the backbone, we use SDXL and fine-tune it using both LoRA and Lily. First, we fine-tune the model on images paired with text prompts (e.g., “A photo of a [v] duck toy”), each of which includes a unique identifier. Afterward, text prompts containing the identifier are used to generate customized images.
Results The results are presented in Fig. 2 following the format in Gao et al. (2024b) and Wu et al. (2024). From these results, we observe that the images generated by Lily generally align better with the text prompts. For instance, when asked to generate an image of a duck toy floating on water, Lily’s output accurately depicts the designated environment, whereas LoRA’s does not. Additionally, when asked to generate an image of a wolf plushie in the snow, Lily precisely captures the snow around the wolf, while LoRA fails to do so. These observations demonstrate Lily’s excellent performance in text-to-image generation with more expressive adaptation. Additional generated results are provided in Appendix I.
@@ 4.4 Visual Adaptation Benchmark
Table 5: Full results of Lily on Vim-S pre-trained on ImageNet-1K for the VTAB-1K benchmark, with averages calculated within each group. * denotes linear probing results from Tu et al. (2023a). For fair comparison, we also use ViT-B pre-trained on ImageNet-1K. Bold indicates best performance among Vim-based PEFT methods.
@@
[…]
Results The results are shown in Tables 4 and 5. For ViT, Lily significantly outperforms all compared PEFT methods while also offering improved parameter efficiency. In contrast, the performance on the Vim backbone is generally lower than that on ViT; for instance, LoRA on ViT performs better than LoRA on Vim. We argue that this difference is due to variations in architecture design and overall model size. However, Lily’s strong adaptation performance allows it to match or exceed the performance of other PEFT methods on ViT and to significantly outperform LoRA on Vim (with both Lily-S and Lily-L surpassing LoRA by a significant margin). This demonstrates Lily’s architecture-agnostic capability, highlighting its potential across various model architectures. Overall, Lily achieves excellent visual adaptation performance while maintaining architecture-agnosticity and high parameter efficiency.
@@ 4.5 Understanding Lily
4.5.1 Does It Have High-Rank Weight Updates?
The interconnected and asymmetric structure of Lily enables a flexible allocation of the parameter budget, thereby allowing weight updates with higher ranks across all layers. To validate this claim, we provide an empirical analysis, as shown in Fig. 3. Specifically, we run four tasks from the NLU experiment and measure the rank of the weight updates for the query transformation matrix $W_{q}$ in the first three layers. We use a small number of matrices $A$ and $B$ (2 or 3) with a rank of 32 to match the parameter count of LoRA, which uses adapters with a rank set to 8. Specific hyperparameter settings can be found in Appendix B.1.2.
From the results, we observe that the rank of the weight updates from Lily is generally notably larger than that of LoRA when using a similar number of parameters. Meanwhile, the weight updates from Lily still exhibit a higher rank compared to those of LoRA even when using only $16.7\%$ of LoRA’s parameters. This empirical analysis essentially validates our claim that Lily achieves high-rank updates with the same parameter budget. We attribute this to the model-wide sharing and the cross-layer asymmetric design, which facilitate a flexible allocation of the parameter budget.
Figure 3: Actual rank of the weight updates. The weight updates are of shape $768\times 768$. We run 20 epochs for COLA, MRPC, and STS-B, and 3 epochs for SST-2. It can be easily observed that the weight updates from Lily have notably higher rank than those from LoRA. Note that the reported rank is computed from accumulated weight updates over multiple epochs.
@@ 4.5.1 Does It Have High-Rank Weight Updates?
Figure 4: Visualization of accumulated assigned weight for $B$ experts by a router across various layers. Example here uses layer of index 2, 13 and 22 to represent shallow, middle and deep layers. The reported values are based on the accumulated router outputs over multiple epochs.
Figure 5: Impact of attention granularity (i.e., the choice of how many $A$s and $B$s) on the performance. We choose 12 out of 19 tasks from VTAB-1K for a comprehensive understanding.
Figure 6: Hardware efficiency of Lily compared to LoRA. We run 10 epochs for COLA. We report the training time and memory consumption. It can be observed that Lily generally performs on par with LoRA in terms of hardware efficiency.
4.5.2 What’s the Influence of Adapter Granularity?
@@
The number of experts in the model-wide $B$ module can be freely set, and the number of $A$s can also be flexibly determined by sharing across the same level of layers, as introduced in Appendix A.1. Therefore, we analyze the impact of these choices on performance. We denote the number of $A$ experts and $B$ experts as ne_1 and ne_2, respectively. For simplicity, we set them equal in the experiments and denote this common value as ne. We refer to the number of layers each expert attends to as adapter granularity. As the value of ne increases, the adapter granularity becomes finer. As shown in Fig. 5, the results from the VTAB-1K benchmark indicate different patterns. For instance, on the DTD dataset, the best performance is achieved when ne is 4, while on sNORB-Azim, performance increases as ne increases. Increasing ne leads to more parameters and finer adapter granularity; however, finer adapter granularity does not necessarily translate to better overall performance. For example, on Resisc45, DTD, Cifar100, sNORB-Ele, dsPr-LoC, Flowers102, and EuroSAT, the negative impact of increasingly finer adapter granularity eventually outweighs the benefits of the additional parameters, leading to a decrease in overall performance. In other tasks, different patterns may occur because the positive effect of adapter granularity on performance is consistently strong, or its negative effect is insufficient to offset the benefits of increased parameters, resulting in generally improved performance with higher ne. This phenomenon provides an important insight: for most tasks, simply increasing the number of parameters may not lead to better performance. Instead, only when adapter granularity and the number of parameters reach an optimal trade-off can we achieve the best performance.
4.5.3 Does It Exhibit Selectivity?